OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -1/4 and passes through the point (8, -1).

OpenStudy (anonymous):

that is easy you already have your slope just fine your y intercept which is where your line crosses y and us y=mx+b or y=1/4x+ y intercept

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem

OpenStudy (karatechopper):

y+1=-1/4(x-8) y+1=-1/4x+4 y+1-1=-1/4x+4-1 y=-1/4x+3

OpenStudy (anonymous):

thank you very much